A tennis ball is traveling at 40 m/s and has a kinetic energy of 80 J. Calculate the mass of the tennis ball. *


1
Expert's answer
2022-02-23T08:30:42-0500

The kinetic energy of a ball

"E_k=\frac{mv^2}{2}"

Hence, the mass of a ball

"m=\frac{2E_k}{v^2}=\frac{2*80}{40^2}=0.1\:\rm kg"